Can you solve this with your work : 1x+2x+3x+4x+5x+6x+..........+98x+99x+100x=50x+10000
P.s. can you show all your work

Answer by stanbon(75887)

Can you solve this with your work : 1x+2x+3x+4x+5x+6x+..........+98x+99x+100x=50x+10000
------
1x+100x = 101x
2x+99x = 101x
3x+98x = 101x
....
...
The left side adds up to (100/2)(1+100)x = 50*101x = 5050x
---------------------
Equation:
5050x = 50x + 10000
5000x = 10,000
x = 2
================
Cheers,
Stan H.